OpenStudy (anonymous):

Solve the system of equations below. x-2y=3 2x-3y=9 A. (3,9) B. (9,3) C. (-3,9) D. (-9,3) ***My answer: B. (9,3) am i right?? thanks!

OpenStudy (anonymous):

hi! thanks for coming :) am i right?

OpenStudy (tamtoan):

if you want to check the answer, just sub your answer in one of the equation and if they come out right, then you got the right answer

OpenStudy (unklerhaukus):

x-2y=3 2x-3y=9 check your answer like this (x,y)=(9,3) 9-2(3)=3 9-6= 3 true 2(9)-3(3)=9 18-9=9 9=9 true

OpenStudy (anonymous):

so i am right @UnkleRhaukus :)

OpenStudy (unklerhaukus):

true and true

OpenStudy (anonymous):

yay thx =)